Set up a system of equations to describe this situation. Solve the system of solutions to find the solution.
A company gets a major part used in the production of their product from two different suppliers. The cost for this major part from dealer A is $5.25 per unit. From supplier B the part cost is $4.80 per unit. The company needs 7500 units and has $37,350 to spend.

Answer by rcmcc(152)

Asuming the company spends $37350 buying the parts we will solve this system in a linear equation.
Let the part costing $5.25 be X and
Let the part costing $4.80 be Y
X+Y=7500 parts
5.25X+4.80Y=37350 spent
multiply the first equation by 5.25 and simplify for y
5.25X+5.25Y=39375
5.25X+4.8Y=37350
0.45Y=2025
y=2025/0.45
y=4500
substitute Y into the first equation to solve for X
X+4500=7500
X=3000
The company buys 3000 of the parts that cost $5.25 and buys 4500 of the parts that cost $4.80